# Proton decay in de Sitter environment

### v1

@@ I Introduction
The Unruh effect has been applied to the problem of the decay of the accelerated proton, $p^{+}\rightarrow n+e^{+}+\nu_{e}$, which is a kind of the inverse $\beta$-decay, see e.g. Refs.Matsas2001 ; Suzuki2002 ; Ahluwalia2016 ; Matsas2017 ; Blasone2020 . Decay of the non-inertial proton was discussed in the earlier papers, see Refs.Ginzburg1965 ; Ginzburg1965b and references therein. Here we consider the similar proton decay in a different configuration – in the de Sitter environment. The de Sitter spacetime has high symmetry due to constant in space and time scalar Riemann curvature. This allows us to study different processes in the de Sitter state using its local thermodynamics.Volovik2024 The de Sitter thermodynamics is characterized by the local temperature $T=H/\pi$, which is twice the Gibbons-Hawking temperature, $T=2T_{\rm GH}$. However, the holographic bulk-surface correspondence does work: the entropy integrated over the Hubble volume coincides with the Gibbons-Hawking entropy $A/4G$, where $A$ is the area of the cosmological horizon.
@@ II.2 General case
Keeping in mind the inverse $\beta$-decay, $p^{+}\rightarrow n+e^{+}+\nu_{e}$, we consider the general case, when the initial particle with mass $M_{\rm in}$ decays into $N$ particles with masses $m_{a}$ such that the total mass $M_{\rm fin}$ of the radiated particles is larger than $M_{\rm in}$:
@@ III.1 Decay of proton in de Sitter environment
$\Gamma(p^{+}\rightarrow n+e^{+}+\nu_{e})\sim\exp{\left(-\frac{\pi\Delta M}{H}\right)}\,,$ (16)
@@ III.2 Proton as the source of electron-positron pairs
If the neutrino mass is neglected, one has
@@
$\Gamma(p^{+}\rightarrow p^{+}+\nu_{e}+\bar{\nu}_{e})\sim\exp{\left(-\frac{2\pi m_{\nu}}{H}\right)}\,.$ (21)

### v2

@@ I Introduction
The Unruh effect has been applied to the problem of the decay of the accelerated proton, $p^{+}\rightarrow n+e^{+}+\nu$, which is a kind of the inverse $\beta$-decay, see e.g. Refs.Matsas2001 ; Suzuki2002 ; Ahluwalia2016 ; Matsas2017 ; Blasone2020 . Decay of the non-inertial proton was discussed in the earlier papers, see Refs.Ginzburg1965 ; Ginzburg1965b and references therein. Here we consider the similar proton decay in a different configuration – in the de Sitter environment. The de Sitter spacetime has high symmetry due to constant in space and time scalar Riemann curvature. This allows us to study different processes in the de Sitter state using its local thermodynamics.Volovik2024 The de Sitter thermodynamics is characterized by the local temperature $T=H/\pi$, which is twice the Gibbons-Hawking temperature, $T=2T_{\rm GH}$. However, the holographic bulk-surface correspondence does work: the entropy integrated over the Hubble volume coincides with the Gibbons-Hawking entropy $A/4G$, where $A$ is the area of the cosmological horizon.
@@ II.2 General case
Keeping in mind the inverse $\beta$-decay, $p^{+}\rightarrow n+e^{+}+\nu$, we consider the general case, when the initial particle with mass $M_{\rm in}$ decays into $N$ particles with masses $m_{a}$ such that the total mass $M_{\rm fin}$ of the radiated particles is larger than $M_{\rm in}$:
@@ III.1 Decay of proton in de Sitter environment
$\Gamma(p^{+}\rightarrow n+e^{+}+\nu)\sim\exp{\left(-\frac{\pi\Delta M}{H}\right)}\,,$ (16)
@@
In a similar way one can consider the following channel of decay of the hydrogen atom:
$\Gamma(H\rightarrow n+\nu)\sim\exp{\left(-\frac{\pi\Delta M}{H}\right)}\,,$ (18)
$\Delta M=M_{\rm fin}-M_{\rm in}=m_{n}+m_{\nu}-m_{H}\,.$ (19)
@@ III.2 Proton as the source of electron-positron pairs
If the neutrino mass is neglected, the mass difference in proton decay in Eq.(17) is:
@@
$\Gamma(p^{+}\rightarrow p^{+}+\nu+\bar{\nu})\sim\exp{\left(-\frac{2\pi m_{\nu}}{H}\right)}\,.$ (23)
@@ IV Discussion
It would be interesting to consider, how these two sides of thermodynamics (the local thermodynamics and the thermodynamics related to the cosmological and Rindler horizons) are modified in the time-dependent environment. The particular case is when the de Sitter state is perturbed towards the Friedmann–Lemaitre–Robertson–Walker cosmology.Odintsov2024